OpenStudy (anonymous):

Which of the following starts with a contradiction of what you are trying to prove? A. contrapositive inference B. direct proof C. indirect proof D. negation

OpenStudy (anonymous):

this doesent have anything to do with math but I think B @amistre64

OpenStudy (freckles):

This is definitely a math question.

OpenStudy (amistre64):

its math, just a subtopic in math

OpenStudy (amistre64):

direct proofs are not by contradiction.

OpenStudy (anonymous):

am I right

OpenStudy (freckles):

Let's think about the option B...If something is direct, do you think that there also a contradiction method involve?

OpenStudy (amistre64):

we can prove 2+2 is equal to 4 by a direct proof. 2+2 = 1+1+1+1 = 4

OpenStudy (anonymous):

would it be d then @amistre64

OpenStudy (amistre64):

its all about definition at the moment. can you define a and d for me?

OpenStudy (anonymous):

ohh i get it now